Database last updated on Tue, 27 Jan 2026 13:40:49 +0000 by localhost

Critical warnings from b-files' comments for days between 2025106 and 2025106 (0 sec, 3 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2025 105 05 32 58 New value set for watermark SUPPLY.VOLTAGE.HIGH343B10625.191TORON#19105:25:221-1
hp:Spectrometer alignment trouble.61B10625.191TORON#19109:20:141-1
re:Previous RE didn11B10625.191TORON#19110:14:031-1

Instrument constant changes for the week before 2025106 (2 sec, 0 rows)

SL burned out? (AP results when SL is ON) (0 sec, 8 rows)

HG burned out? (AP results when HG is ON) (0 sec, 11 rows)

Non-critical warnings from b-files' comments for days between 2025106 and 2025106 (0 sec, 5 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2025 105 05 31 56 Reset Warm -> Tepid from 2025 105 05 31 50 until 2025 105 05 31 51 . Counts: 1,013B10625.191TORON#19105:25:191-1
rl:2025 106 10 12 16 Reset Warm -> Warm from 2025 106 10 12 13 until 2025 106 10 12 14 . Counts: 0,123B10625.191TORON#19105:26:171-1
re:Zenith discrepancy = -11399B10625.191TORON#19105:33:13-11-1010
pd:Going to menu199B10625.191TORON#19110:14:051-1
ds:Suppressed oscillation of filter wheel.35B10625.191TORON#19120:25:491-1

Measurements and tests summary.


_jday_loc_serial_skcaodedsdtedfmfzhghprsslzs
106TORONTO#191ecnouv,Next schedule will be ecnouv (now ecnouv ZS= 93 ZS= 93.386)89942211117242742
(1 sec)

Relative and absolute humidity for day 2025106 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
19110.31.6